OpenStudy (anonymous):

Solve: 5(x + 8) = 7(x − 2)

OpenStudy (anonymous):

\[\Large \begin{array}{l} 5(x + 8) = 7(x - 2)\\ 5x + 40 = 7x - 14\\ - 2x = - 54\\ x = \frac{{54}}{2} = 27 \end{array}\]

OpenStudy (anonymous):

you need to simplify it first, so you use distributive property. Solve: 5(x + 8) = 7(x − 2) = 5x + 40 = 7x -14 then you move all the variables (x) onto one side and all the other numbers (+40 and -14) onto the other side. 5x +40 (-40) = 7x -14 (-40) 5x= 7x -54 5x (-7x) = 7x (-7x) -54 -2x = -54 then you isolate for x... (-2x ÷ -2) = (-54 ÷-2) x= -27 :)